# Hear What Matters! Text-conditioned Selective Video-to-Audio Generation

### v1

@@ 4 Experiments
ReWaS [rewas] 70.4 4.937  6.23 2.57 0.200 0.2454 1.364
VinTAGe [vintage] 50.5 1.309 11.51 1.69 0.283 0.2850 1.292
MMAudio-S-16k [mmaudio] 56.7 0.874 11.54 2.07 0.270 0.3135 0.802
VOS [deva]+MMAudio [mmaudio] 60.0 0.878 12.11 1.91 0.291 0.3010 0.991
@@
ReWaS [rewas] 57.4 3.148 6.29 1.97 0.220 0.2569 1.377
VinTAGe [vintage] 37.0 0.690 9.28 0.88 0.277 0.2892 1.304
MMAudio-S-16k [mmaudio] 41.5 0.654 9.00 1.09 0.276 0.3248 0.670
VOS [deva]+MMAudio [mmaudio] 43.4 0.656 8.91 1.11 0.287 0.3087 0.904
@@ 4.1 Setup
SelVA is trained on VGGSound [vggsound], which provides approximately 500 hours of video and 310 unique captions. We utilize these captions as the text prompts for our model. Following the experimental setup of our baseline model [mmaudio], we partition the official training data, setting aside 2k samples for validation. This results in a 179k-sample training set, a 2k-sample validation set, and the official 15k-sample test set. For both training and inference, all video clips are processed into 8-second segments.
@@
Evaluating selective sound generation requires clean, source-separated audio with corresponding text descriptions. However, existing in-the-wild datasets such as VGGSound [vggsound] and AudioSet [audioset] typically provide only a single mixed track and video-level captions, often contaminated by recording noise or off-screen sounds [regnet, varietysound, vintage]. To address these limitations, we introduce VGG-MonoAudio, an evaluation benchmark for selective V2A generation. We curate mono-source clips from UnAV-100 [unav100] overlapping with VGGSound test set, and filter them automatically and manually with three strict criteria: (1) a single source sounding with minimal background or off-screen noise, (2) the sounding object is clearly visible, (3) the text annotation precisely matches the auditory event. Finally, we obtain a final set of 67 clean, single-source videos spanning 39 unique events (e.g., ‘dog barking’, ’train wheels squealing’) across 8 categories: human, music, vehicle, tool, animal, nature, sport, other. To construct test samples, we concatenate pairs of these videos side-by-side, each occupying half the frame width. The horizontally combined video $V_{1}\oplus V_{2}]$ serves as the model input, while the audio $A_{\mathbf{v}_{1}}$ and text $T_{\mathbf{v}_{1}}$ from one of the source videos are used as the target. From the 67 curated videos, we generate 1,071 mixed pairs in total, 560 inter-class (videos from different categories) and 511 intra-class (videos from the same category), for quantitative evaluation. Appendix B provides the full list of audio event categories and detailed benchmark statistics. We also include results on the original VGGSound test set for completeness; these are given in Appendix D.2, as this setting is not central to our evaluation.
@@
We establish four SoTA baselines for comparison. ReWaS [rewas], VinTAGe [vintage], and MMAudio [mmaudio] are text-conditioned V2A models where text semantically aids video. To implement concurrent segmentation-based approaches [hearyourclick, saganet], we leverage the pretrained video object segmentation (VOS) model (i.e., DEVA [deva]) to build a ‘VOS+MMAudio’ system, in which we first pass the text prompt and video to obtain video-level segmentation masks. The resulting masked video is then used as the conditional input to the MMAudio.
@@
  • Audio quality: Overall sound quality is assessed by Fréchet audio distance (FAD) [fad], kernel audio distance (KAD) [kad], and inception score (IS) [inception_score] with PANNs[panns, correlation_fad].
  • Semantic alignment: Semantic similarities are assessed to evaluate prompt fidelity. While CLAP score (CLAP) [clap_laion] is used to measure how closely the generated audio aligns with the intended text, imagebind score (IB) [imagebind] measures between audio and target video. In addition, Kullback-Leibler divergence (KL) with PANNs distribution is employed to evaluate semantic alignment between the generated and groundtruth audio tracks.
  • Temporal alignment: We report DeSync [mmaudio], the average synchronized error (i.e., predicted offset in seconds) between the audio and video. As temporal alignment is crucial for perceptual coherence in V2A, this metric serves as the primary reference in our ablation study.
@@ 4.2 Implementation details
During training of SelVA, mixing inputs are given within each minibatch with a probability of 0.75, while clipping the mixing ratio $\lambda$ of the target video to be greater than 0.2. A total of 5 learnable [SUP] tokens are prepended to every input text prompt; this number was determined by Tab. D3. We initialize the video encoder $\mathcal{F}_{S}$ with pretrained Synchformer [synchformer] and the generator $\mathcal{G}$ with MMAudio-small-16kHz weights. Note that we train 19M parameters in $\mathcal{F}_{S}$ and 22M in $\mathcal{G}$, corresponding to 14% of each model’s total parameters, respectively. Following the original setup for classifier-free guidance (CFG) in MMAudio [mmaudio], we randomly substitute the video and text features with learned null video and text embedding ($\varnothing_{\mathbf{v}}$ and $\varnothing_{\mathbf{t}}$) with a probability of 0.1. In addition, we drop the text feature with an additional probability of 0.5 to enhance the visual fidelity. Inference on the flow matching model is performed using the Euler solver with 25 linear sampling steps. During inference, CFG is applied with a guidance strength of $\gamma=4.5$.
@@ 4.3 Comparison with state-of-the-arts
Table 1 summarizes the quantitative performance of V2A models on VGG-MonoAudio. SelVA outperforms baselines across all key aspects, including audio quality, semantic alignment, and temporal alignment. Notably, we achieve the best scores in both semantic and temporal audio-video alignment. MMAudio [mmaudio], which overlooks text modality, exhibits degraded CLAP scores than SelVA, whereas neglecting video modality often results in temporally misaligned results with poor DeSync scores, as seen in ReWaS [rewas] and VinTAGe [vintage]. This highlights that training a text-conditioned video encoder in SelVA is effective to achieve these dual goals. VOS baseline shows competitive semantic alignment but performs poorly on temporal synchronization. It is primarily due to the inherent limitations of VOS methods, which struggle to accurately localize fast-moving or motion-blurred objects, and vague or complex boundaries (e.g., rain drop). Regarding the two subsets of VGG-MonoAudio, models generally achieve better objective scores in the intra-class subset. This happens because the paired non-target video is semantically similar to the target, leading objective metrics to overestimate model performance. Therefore, human perceptual evaluation becomes essential.
@@
Fig. 4 demonstrates that SelVA selectively synthesizes target sounds even in complex real-world auditory scenes involving multiple simultaneous or temporally overlapping events. As shown in Fig. 4(a), SelVA successfully generates distinct sounds for spatially co-occurring sound events such as ‘car passing by’ and ‘man running’, demonstrating spatial disentanglement within a shared visual context. In Fig. 4(b), SelVA succeeds in temporally disentangling intertwined events, producing natural and temporally aligned sounds for ‘firework banging’ and ‘people speaking’. These examples highlight our model’s robustness to yield realistic and context-aware audio by capturing the user’s intended focus. More qualitative examples are provided in Fig. D7 in Appendix, and audio samples are available in the supplementary material.
@@ 4.4 Ablation studies
Achieving the dual goal of semantic text-audio alignment and temporal video-audio alignment is crucial in selective sound generation. We therefore observe the change in both the CLAP and DeSync scores for different numbers of [SUP] tokens. As shown in Fig. 6, we identify a “sweet spot” at 5 tokens, which achieves the best DeSync score while maintaining a comparable CLAP score. It is a consistent observation from prefix-tuning [prefix_tuning], where performance typically degrades when too few tokens fail to convey sufficient conditioning information, or when too many tokens lead to redundancy and overfitting. Full results, including other metrics, are shown in Tab. D3 in Appendix.
@@ 5 Limitations and Future Work
We identify three primary directions for future work. First, the model performance is currently limited by the noisiness of the training data in VGGSound [vggsound]. Therefore, more rigorous data filtering or refining the auto-mixing process with cleaner source data could improve the performance. Second, since text labels are typically simple noun-verb conjunctions and lack such descriptive detail, the model’s complex text understanding capabilities could be enhanced. This includes fine-grained cross-modal distinction (e.g., separating ‘male singing’ from ‘male burping’) and improved attribute controllability (e.g., a dog barking ‘aggressively’). Finally, while our method significantly alleviates the sound substitution issue, residual cases remain when the video encoder fails to track a target movement change consistently. We leave a comprehensive full training of the model as future work.
@@ 6 Conclusion
We present SelVA, text-conditioned V2A model tailored for audio production systems in the real world. SelVA efficiently modulates the video encoder to capture the user’s textual intent, introducing a few learnable tokens and specialized training schemes. Experimental results show that SelVA delivers precise and controllable sound generation on our new benchmark, VGG-MonoAudio, significantly outperforming existing methods. These findings highlight SelVA as a strong step toward practical, reliable, and fully controllable V2A generation.
[…]
Supplementary Material
@@ Appendix A More Implementation Details
Table A1: Performance of MMAudio [mmaudio] on VGGSound test set with different input visual feature combinations.
@@ A.1 Video encoder
We first analyze the role of each vision feature (i.e., CLIP [clip] and Synchformer [synchformer]) used in MMAudio [mmaudio], which is also used in parameter initialization of our method. Understanding how each embedding contributes to generation quality is crucial for determining how to adapt the video encoder for text conditioning. Tab. A1 reports an ablation study by substituting each feature with its corresponding learned null embedding trained for classifier-free guidance. Interestingly, removing the CLIP [clip] embedding actually improves the CLAP score by 0.038 while the DeSync score remains steady. This suggests that the CLIP embedding often introduces semantic distraction without conveying significant temporal information. In contrast, the IB and DeSync scores deteriorate significantly when the Synchformer embedding is removed. This shows that the Synchformer feature contributes both semantic and temporal information for reliable audio-video alignment.
@@
Synchformer, consisting of audio encoder and video encoder, is learned to predict the temporal offset to evaluate audiovisual synchronization. In this experiment, we only use the video encoder for feature extraction, following the details of MMAudio. Note that the architecture of the video encoder follows the Motionformer with divided space-time attention [motionformer, timesformer]. Given an input video of 8 seconds at 25 fps, we first divide it into segments with windowing (window size of 16, hop size of 8 frames), which results in 24 segments. Here, a batched video data has a shape [Batch, Segments, Channel, Height, Width] by resizing $224\times 224$ resolution without center crop. Each video frame is patchified and flattened in rasterized order. After passing the video encoder, each segment results in $8$ embeddings in the temporal axis with a hidden dimension of $D=768$. The final video feature $\mathbf{v}$ of a minibatch has a shape of [Batch, Segments, $t=8$, $D=768$].
@@ A.2 Text encoder
To extract text embeddings, we use Flan-T5-Base [flant5]  [^1] https://huggingface.co/google/flan-t5-base to condition the video encoder. For the audio generator, to reuse pretrained parameters from MMAudio, CLIP’s text encoder [clip] is employed.
@@ A.3 Training
In the first training stage, we finetune the Synchformer [synchformer] video encoder to condition text prompts. We use the pretrained checkpoint 24-01-04T16-39-21 from the official implementation [^2] https://github.com/v-iashin/Synchformer, trained on AudioSet [audioset] using a two-stage process consisting of audio-visual contrastive learning and offset estimation. Specifically, we train the initialized spatial attention pooling layer and a new trainable text cross-attention block, which is placed after the space-time attention blocks (see Fig. A1 for details). This process ensures parameter-efficient finetuning, updating only 14% (19M) of the 135M total parameters. We train for 50k steps with a batch size of 4 on a single NVIDIA RTX 4090. The base learning rate is set to 1e-4, with a 1k-step linear warmup schedule.
In the second training stage, we train the multimodal-conditioned audio generator. To efficiently train the large-scale generator, we take the initial parameters from the MMAudio-small-16k model [mmaudio]. Therefore, the architecture of the generator in SelVA is similar to MMAudio. Only 14% (22M) of the 157M total parameters are trained in our second stage. Concretely, we finetune the initial projection layer for the Synchformer video feature $\mathbf{v}_{\text{Sync}}$ and all adaLN-related layers that receive the video feature as input. Fig. A2 specifies those learnable layers within a single MM-DiT [sd3, flux] block of MMAudio. It is worth noting that a frame-aligned conditioning $c_{f}$ is a function of the Synchformer video feature, while the global conditioning $c_{g}$ is not. While MMAudio originally used the CLIP image feature $\mathbf{v}_{\text{CLIP}}$, we do not use this for conditioning by replacing with the null feature $\varnothing_{\mathbf{v}}$. We train for 25k steps with a batch size of 12 on a single NVIDIA RTX A6000. The base learning rate is set to 1e-5, also with a 1k-step linear warmup.
Common to both training stages, we utilize bf16 automatic mixed precision (AMP). We employ the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 1e-6, along with gradient clipping at a norm of 1. After training, we apply post-hoc EMA [posthoc_ema] with a relative width of $\sigma_{\text{rel}}=0.05$.

### v2

@@ 4 Experiments
ReWaS [32] 70.4 4.937  6.23 2.57 0.200 0.2454 1.364
VinTAGe [39] 50.5 1.309 11.51 1.69 0.283 0.2850 1.292
MMAudio-S-16k [8] 56.7 0.874 11.54 2.07 0.270 0.3135 0.802
VOS [9]+MMAudio [8] 60.0 0.878 12.11 1.91 0.291 0.3010 0.991
@@
ReWaS [32] 57.4 3.148 6.29 1.97 0.220 0.2569 1.377
VinTAGe [39] 37.0 0.690 9.28 0.88 0.277 0.2892 1.304
MMAudio-S-16k [8] 41.5 0.654 9.00 1.09 0.276 0.3248 0.670
VOS [9]+MMAudio [8] 43.4 0.656 8.91 1.11 0.287 0.3087 0.904
@@ 4.1 Setup
SelVA is trained on VGGSound [5], which provides approximately 500 hours of video and 310 unique captions. We utilize these captions as the text prompts for our model. Following the experimental setup of our baseline model [8], we partition the official training data, setting aside 2k samples for validation. The training set includes 179k videos, and the test set is 15k. For both training and inference, all video clips are processed into 8-second segments.
@@
Evaluating selective sound generation requires clean, source-separated audio with corresponding text descriptions. However, existing in-the-wild datasets such as VGGSound [5] and AudioSet [21] typically provide only a single mixed track and video-level captions, often contaminated by recording noise or off-screen sounds [6, 13, 39]. To address these limitations, we introduce VGG-MonoAudio, an evaluation benchmark for selective V2A generation. We curate mono-source clips from UnAV-100 [22] overlapping with VGGSound test set, and filter them automatically and manually with three strict criteria: (1) a single source sounding with minimal background or off-screen noise, (2) the sounding object is clearly visible, (3) the text annotation precisely matches the auditory event. Finally, we obtain a final set of 67 clean, single-source videos spanning 39 unique events (e.g., ‘dog barking’, ’train wheels squealing’) across 8 categories: human, music, vehicle, tool, animal, nature, sport, other. To construct test samples, we concatenate pairs of these videos side-by-side, each occupying half the frame width. The horizontally combined video $V_{1}\oplus V_{2}]$ serves as the model input, while the audio $A_{\mathbf{v}_{1}}$ and text $T_{\mathbf{v}_{1}}$ from one of the source videos are used as the target. From the 67 curated videos, we generate 1,071 mixed pairs in total, 560 inter-class (videos from different categories) and 511 intra-class (videos from the same category), for quantitative evaluation. Appendix B provides the full list of audio event categories and detailed benchmark statistics. We also include results on the original VGGSound test set for completeness; these are given in Appendix D.2, as this setting is not central to our evaluation.
@@
We establish four SoTA baselines for comparison. ReWaS [32], VinTAGe [39], and MMAudio [8] are text-conditioned V2A models where text semantically aids video. To implement concurrent segmentation-based approaches [45, 60], we leverage the pretrained video object segmentation (VOS) model [9] to build a ‘VOS+MMAudio’ system, in which we pass the text and video to obtain video-level object masks. The resulting masked video is then used as the conditional input to the MMAudio.
@@
  • Audio quality is assessed by Fréchet audio distance (FAD) [34], kernel audio distance (KAD) [11], and inception score (IS) [56] with PANNs[37, 58].
  • Semantic alignment is assessed to evaluate prompt fidelity. While CLAP score (CLAP) [65] is used to measure how closely the generated audio aligns with the intended text, imagebind score (IB) [23] measures the alignment between audio and target video. In addition, Kullback-Leibler divergence (KL) with PANNs distribution is employed to evaluate semantic alignment between the generated and groundtruth audio tracks.
  • Temporal alignment is assessed by DeSync [8], the average synchronized error (i.e., predicted offset in seconds) between the audio and video. As temporal alignment is crucial for perceptual coherence in V2A, this metric serves as the primary reference in our ablation study.
@@ 4.2 Implementation details
During training of SelVA, mixing inputs are given within each minibatch with a probability of 0.75, while clipping the mixing ratio $\lambda$ of the target video to be greater than 0.2. A total of 5 learnable [SUP] tokens are prepended to every input text prompt; this number was determined by Tab. D3. We initialize the video encoder $\mathcal{F}_{S}$ with pretrained Synchformer [31] and the generator $\mathcal{G}$ with MMAudio-small-16kHz weights. Note that we train 19M parameters in $\mathcal{F}_{S}$ and 22M in $\mathcal{G}$, corresponding to 14% of each model’s total parameters, respectively. Following the original setup for classifier-free guidance (CFG) in MMAudio [8], we randomly substitute the video and text features with learned null video and text embedding ($\varnothing_{\mathbf{v}}$ and $\varnothing_{\mathbf{t}}$) with a probability of 0.1. In addition, we drop the text feature with an additional probability of 0.5 to enhance the visual fidelity. Inference on the flow matching model is performed using the Euler solver with 25 linear sampling steps. During inference, CFG is applied with a guidance strength of $\gamma=4.5$.
@@ 4.3 Comparison with state-of-the-arts
Table 1 summarizes the quantitative performance of V2A models on VGG-MonoAudio. SelVA outperforms baselines across all key aspects, including audio quality, semantic alignment, and temporal alignment. Notably, we achieve the best scores in both semantic and temporal audio-video alignment. MMAudio [8], which overlooks text modality, exhibits degraded CLAP scores than SelVA, whereas neglecting video modality often results in temporally misaligned results with poor DeSync scores, as seen in ReWaS [32] and VinTAGe [39]. This highlights that training a text-conditioned video encoder in SelVA is effective to achieve these dual goals. VOS baseline shows competitive semantic alignment but performs poorly on temporal synchronization. It is primarily due to the inherent limitations of VOS methods, which struggle to accurately localize fast-moving or motion-blurred objects, and vague or complex boundaries (e.g., rain drop). Regarding the two subsets of VGG-MonoAudio, models generally achieve better objective scores in the intra-class subset. This happens because the paired non-target video is semantically similar to the target, leading objective metrics to overestimate model performance. Therefore, human perceptual evaluation becomes essential.
@@
Fig. 4 demonstrates that SelVA selectively synthesizes target sounds even in complex real-world auditory scenes involving multiple simultaneous or temporally overlapping events. As shown in Fig. 4(a), SelVA successfully generates distinct sounds for spatially co-occurring sound events such as ‘car passing by’ and ‘man running’, demonstrating spatial disentanglement within a shared visual context. In Fig. 4(b), SelVA succeeds in temporally disentangling intertwined events, producing natural and temporally aligned sounds for ‘firework banging’ and ‘people speaking’. These examples highlight our model’s robustness to yield realistic and context-aware audio by capturing the user’s intended focus. More qualitative examples are provided in Fig. D8 in Appendix.
@@ 4.4 Ablation studies
Achieving the dual goal of semantic text-audio alignment and temporal video-audio alignment is crucial in selective sound generation. We therefore observe the change in both the CLAP and DeSync scores for different numbers of [SUP] tokens. As shown in Fig. 6, we identify a “sweet spot” at 5 tokens, which achieves the best DeSync score while maintaining a comparable CLAP score. It is a consistent observation from prefix-tuning [44], where performance typically degrades when too few tokens fail to convey sufficient conditioning information, or when too many tokens lead to redundancy and overfitting. Full results, including other metrics, are shown in Tab. D3 in Appendix.
@@ 5 Limitations and Future Work
We identify three primary directions for future work. First, the model performance is currently limited by the noisiness of the training data in VGGSound [5]. Therefore, more rigorous data filtering or refining the auto-mixing process with cleaner source data could improve the performance. Second, since text labels are typically simple noun-verb conjunctions and lack such descriptive detail, the model’s complex text understanding capabilities could be enhanced. This includes fine-grained cross-modal distinction (e.g., separating ‘male singing’ from ‘male burping’) and improved attribute controllability (e.g., a dog barking ‘aggressively’). Finally, while our method significantly alleviates the sound substitution issue, residual cases remain when the video encoder fails to track a target movement change consistently. We leave a comprehensive full training of the model as future work.
@@ 6 Conclusion
We present SelVA, text-conditioned V2A model tailored for audio production systems in the real world. SelVA efficiently modulates the video encoder to capture the user’s textual intent, introducing a few learnable tokens and specialized training schemes. Experimental results show that SelVA delivers precise and controllable sound generation on our new benchmark, VGG-MonoAudio, significantly outperforming existing methods. These findings highlight SelVA as a strong step toward practical, reliable, and fully controllable selective video-to-audio generation.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00222383, RS-2025-16065706).
Appendix
@@ Appendix A More Implementation Details
Table A1: Performance of MMAudio [8] on VGGSound test set with different input visual feature combinations.
@@ A.1 Video encoder
We first analyze the role of each vision feature (i.e., CLIP [54] and Synchformer [31]) used in MMAudio [8], which is also used in parameter initialization of our method. Understanding how each embedding contributes to generation quality is crucial for determining how to adapt the video encoder for text conditioning. Tab. A1 reports an ablation study by substituting each feature with its corresponding learned null embedding trained for classifier-free guidance. Interestingly, removing the CLIP [54] embedding actually improves the CLAP score by 0.038 while the DeSync score remains steady. This suggests that the CLIP embedding often introduces semantic distraction without conveying significant temporal information. In contrast, the IB and DeSync scores deteriorate significantly when the Synchformer embedding is removed. This shows that the Synchformer feature contributes both semantic and temporal information for reliable audio-video alignment.
@@
Synchformer, consisting of audio encoder and video encoder, is learned to predict the temporal offset to evaluate audiovisual synchronization. In this experiment, we only use the video encoder for feature extraction, following the details of MMAudio. Note that the architecture of the video encoder follows the Motionformer with divided space-time attention [51, 4]. Given an input video of 8 seconds at 25 fps, we first divide it into segments with windowing (window size of 16, hop size of 8 frames), which results in 24 segments. Here, a batched video data has a shape [Batch, Segments, Channel, Height, Width] by resizing $224\times 224$ resolution without center crop. Each video frame is patchified and flattened in rasterized order. After passing the video encoder, each segment results in $8$ embeddings in the temporal axis with a hidden dimension of $D=768$. The final video feature $\mathbf{v}$ of a minibatch has a shape of [Batch, Segments, $t=8$, $D=768$].
@@ A.2 Text encoder
To extract text embeddings, we use Flan-T5-Base [10]  [^1] https://huggingface.co/google/flan-t5-base to condition the video encoder. For the audio generator, to reuse pretrained parameters from MMAudio, CLIP’s text encoder [54] is employed.
@@ A.3 Training
In the first training stage, we finetune the Synchformer [31] video encoder to condition text prompts. We use the pretrained checkpoint 24-01-04T16-39-21 from the official implementation  [^2] https://github.com/v-iashin/Synchformer, trained on AudioSet [21] using a two-stage process consisting of audio-visual contrastive learning and offset estimation. Specifically, we train the initialized spatial attention pooling layer and a new trainable text cross-attention block, which is placed after the space-time attention blocks (see Fig. A1 for details). This process ensures parameter-efficient finetuning, updating only 14% (19M) of the 135M total parameters. We train for 50k steps with a batch size of 4 on a single NVIDIA RTX 4090. The base learning rate is set to 1e-4, with a 1k-step linear warmup schedule.
In the second training stage, we train the multimodal-conditioned audio generator. To efficiently train the large-scale generator, we take the initial parameters from the MMAudio-small-16k model [8]. Therefore, the architecture of the generator in SelVA is similar to MMAudio. Only 14% (22M) of the 157M total parameters are trained in our second stage. Concretely, we finetune the initial projection layer for the Synchformer video feature $\mathbf{v}_{\text{Sync}}$ and all adaLN-related layers that receive the video feature as input. Fig. A2 specifies those learnable layers within a single MM-DiT [18, 40] block of MMAudio. It is worth noting that a frame-aligned conditioning $c_{f}$ is a function of the Synchformer video feature, while the global conditioning $c_{g}$ is not. While MMAudio originally used the CLIP image feature $\mathbf{v}_{\text{CLIP}}$, we do not use this for conditioning by replacing with the null feature $\varnothing_{\mathbf{v}}$. We train for 25k steps with a batch size of 12 on a single NVIDIA RTX A6000. The base learning rate is set to 1e-5, also with a 1k-step linear warmup.
Common to both training stages, we utilize bf16 automatic mixed precision (AMP). We employ the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 1e-6, along with gradient clipping at a norm of 1. After training, we apply post-hoc EMA [33] with a relative width of $\sigma_{\text{rel}}=0.05$.